THE WIDTH OF A RECTANGLE IS 3 FEET LESS THAN ITS LENGHTH .IF THE PERIMETER IS 22 FEET WHAT IS THE WIDTH?: THE WIDTH OF A RECTANGLE IS 3 FEET LESS THAN ITS LENGHTH .IF THE PERIMETER IS 22 FEET WHAT IS THE WIDTH?
Answer by ankor@dixie-net.com(3949)

THE WIDTH OF A RECTANGLE IS 3 FEET LESS THAN ITS LENGHTH .IF THE PERIMETER IS 22 FEET WHAT IS THE WIDTH?
:
Let x = the width
It says the length is 3 ft longer, therefore:
(x+3) = the length
:
Perimeter is 2L + 2W, therefore:
2(x+3) + 2x = 22
:
2x + 6 + 2x = 22
:
2x + 2x = 22 - 6
:
4x = 16
x = 16/4
x = 4 ft is the width
and
4 + 3 = 7 ft is the length
:
:
Check using the given perimeter
2(7) + 2(4) = 22